# Coward: Collision-based Watermark for Proactive Federated Backdoor Detection

### v3

@@ Abstract
Backdoor detection is currently the mainstream defense against backdoor attacks in federated learning (FL), where a small number of malicious clients can upload poisoned updates to compromise the federated global model. Existing backdoor detection techniques fall into two categories, passive and proactive, depending on whether the server proactively intervenes in the training process. However, both of them have inherent limitations in practice: passive detection methods are disrupted by common non-i.i.d. data distributions and random participation of FL clients, whereas current proactive detection methods are misled by an inevitable out-of-distribution (OOD) bias because they rely on backdoor coexistence effects. To address these issues, we introduce a novel proactive detection method dubbed Coward, inspired by our discovery of multi-backdoor collision effects, in which consecutively planted, distinct backdoors significantly suppress earlier ones. Correspondingly, we modify the federated global model by injecting a carefully designed backdoor-collided watermark, implemented via regulated dual-mapping learning on OOD data. This design not only enables an inverted detection paradigm compared to existing proactive methods, thereby naturally counteracting the adverse impact of OOD prediction bias, but also introduces a low-disruptive training intervention that inherently limits the strength of OOD bias, leading to significantly fewer misjudgments. Extensive experiments on benchmark datasets show that Coward achieves state-of-the-art detection performance, effectively alleviates OOD prediction bias, and remains robust against potential adaptive attacks. The code for our method is available at https://github.com/still2009/cowardFL.
Index Terms:
Backdoor Defense, Backdoor Attack, Federated Learning, AI Security, Trustworthy ML.
@@ 1 Introduction
Federated Learning (FL) [mcmahan2017fedavg, fl_big_survey, fl_survey24_pami] has emerged as a powerful paradigm for privacy-preserving machine learning, enabling collaborative model training across distributed clients without sharing raw data. By keeping data locally owned and processed, FL inherently strengthens privacy and copyright protection, and unlocks the value of proprietary, professional, domain-specific and personal data held by private organizations or individuals. This privacy-native property makes FL particularly suitable for securely building personalized and domain-specialized intelligence that goes beyond models trained solely on generic public data [tdsc25_fl_personalize, fl_fm_nips, fan2025_fl_foudationmodel]. As a result, FL has been extensively studied across a wide range of application domains, including healthcare, finance, and edge intelligence [fl_nature_medi24, fl_finance_icse24, fl_edge_survey24].
@@
Figure 1: Our backdoor-collided detection versus the existing backdoor-coexistent method. By developing a backdoor-collided watermark, our method enables an inverted detection paradigm that effectively mitigates misleading effects caused by the OOD prediction bias, a key challenge that limits current proactive detection methods.
@@
However, the privacy-preserving design of federated learning is a double-edged sword. While keeping raw data local effectively protects sensitive information, it also restricts the server’s ability to inspect participant behavior during training. This restricted observability creates an inevitable vulnerability to model-level attacks, where the integrity of the federated global model can be compromised by covertly manipulated local models, most prominently through backdoor attacks [ba_survey_yiming, chameleon, yang2024_ba_atk]. In general, backdoor attacks cause models to behave normally on benign inputs while producing harmful outputs only under specific trigger conditions, rendering the malicious behavior stealthy and difficult to detect. Once the malicious backdoor models are aggregated together with benign ones, the backdoor is seamlessly propagated to the global model, thereby manipulating the entire federated system [bagdasaryan2020backdoor, tdsc25_fba_atk1_durable, tdsc25_fba_atk2_collusive].
@@
To detect and exclude malicious backdoor models, passive detection methods have been proposed [rieger2022deepsight, nguyen2022flame], which do not intervene in the FL process and assume that malicious models manifest observable model-level deviations as outliers. They rely on similarity-based measurements over either model parameters or outputs and identify anomalous models either through direct thresholding or by further applying clustering to expose structurally isolated outliers. However, we reveal that these methods are vulnerable to client non-i.i.d. data distributions and participation randomness, which are common settings in FL scenarios [fl_survey24_pami]. These factors lead to highly dynamic and divergent benign local models that may appear as outliers, significantly interfering with detection. We refer to this limitation as the non-i.i.d. divergence challenge, as analyzed in Section 3.2.
@@
Recently, the pioneering work BackdoorIndicator [li2024backdoorindicator] proposed the first federated proactive detection strategy, which explicitly intervenes in the training process to induce malicious models to exhibit behaviors distinct from benign ones. Specifically, BackdoorIndicator injects OOD patterns designed to be easily retained by malicious client models while tending to be forgotten by benign ones, such that stronger pattern retention in a local model indicates a malicious client. By shifting the detection basis from model discrepancies to differences in client responses to intervention actions, proactive methods naturally avoid the sensitivity to non-i.i.d. distributions that often undermines passive detection methods. However, due to the privacy constraints of FL that prevent access to in-distribution client data, proactive training interventions at the server are restricted to operating via OOD data. As a result, detection outcomes become susceptible to the OOD prediction bias of deep neural networks (DNNs) [nguyen2015ood_bias, guo2017ood_bias_calib, hein2019relu_ood_bias_theory]. This bias causes benign local models to be easily misidentified as malicious clients. We provide a detailed analysis in Section 3.3 and refer to this limitation as the OOD bias challenge.
@@
To tackle above challenges, we propose a simple yet effective Collision-based OOD WAtermark for Robust federated backdoor Detection method, termed Coward. Our method is inspired by the finding of multi-backdoor collision effect where a consecutively planted distinct backdoor will significantly interfere with the one previously planted. To this end, we detect backdoor models by evaluating whether the server-injected conflicting global watermark has been erased during local training. Specifically, we implant a conflicting OOD watermark, where OOD samples embedded with a specific watermark trigger are redirected from their original class to a designated target class. Concretely, the OOD watermark is injected by jointly learning: (1) a regulated base OOD mapping that specifies a fixed class-wise correspondence from OOD samples to the task label space; and (2) a targeted watermark mapping that enforces an association between all triggered OOD samples and the watermark target class. In general, our method alleviates the OOD bias challenge through three aspects: (1) Unlike prior approaches, our collision-based detection treats high inspection accuracy as the signal of benign behavior, making it naturally compatible with the OOD bias–induced high confidence predictions; (2) Our watermark-based detection inspects the presence of the trigger-to-target mapping rather than the OOD sample-to-target mapping, mitigating the direct impact of OOD prediction bias. (3) Our method enforces explicit class-wise label assignments to OOD samples to reduce random prediction fluctuations. We demonstrate (1) and (2) in Figure 1, and verify (3) in Figure 7.
In summary, our contributions are threefold:
  • We revisit existing detection methods for FL backdoor attacks and identify two key challenges inherent to FL settings: non-i.i.d. divergence and OOD prediction bias. We show how these challenges undermine existing detection methods, leading to high benign-client misclassification.
@@
  • We find an intriguing phenomenon, dubbed the multi-backdoor collision effect, where a consecutively planted distinct backdoor can significantly suppress the previously implanted one. Based on this finding, we design a simple yet effective proactive backdoor detection method, $i.e.$, Coward, that effectively mitigates OOD bias and achieves more accurate detection.
  • Extensive experiments on benchmark datasets verify the effectiveness of our method and its resistance to potential adaptive backdoor attacks.
@@ 2.1 Federated Learning
Federated learning enables multiple decentralized clients to collaboratively train a shared model without exchanging their private data [fl_survey24_pami]. Based on how data is partitioned, FL is typically categorized into horizontal FL (HFL), where clients share the same feature space but different samples, and vertical FL (VFL), where clients share samples but differ in input feature fields [flsurvey_yangqiang].
In this paper, we focus on the HFL setting, where a central server coordinates model training by periodically selecting a subset of clients to receive the current global model, update it using local data, and return the updated parameters. The server then aggregates these updates to form a new global model, and this process repeats iteratively, progressively improving model performance while keeping all training data local. Specifically, given the selected client set $\mathcal{S}^{t-1}$ in round t1t-1, the global model $\bm{w}^{t}$ in round $t$ is:
@@
where $N_{k}$ and $N$ denote the data volume of client $k$ and the total data volume across participating clients, respectively. $\Delta$ denotes the model residual with respect to the previous global model. We highlight three key characteristics that distinguish FL from centralized settings:
  • Multi-party Participation: The global model is collaboratively trained by many clients, who privately contributes to the final model but can’t control it individually.
  • Non-i.i.d. Client Data [mcmahan2017fedavg]: Data distributions vary significantly across clients, resulting in inconsistent local updates and increased optimization difficulty.
  • Partial Client Participation [mcmahan2017fedavg]: In each training round, only a subset of clients is randomly selected due to limited communication resources and dynamic client availability.
@@ 2.2 Backdoor Attacks in Federated Learning
Backdoor attacks aim to implant hidden behaviors into neural networks, causing them to function normally on clean data but misclassify inputs containing attacker-specified triggers [ba_survey_yiming, gu2020badnets]. Data poisoning is the most common backdoor injection method, where a subset of training samples is embedded with triggers and mislabeled to a target class [gu2020badnets]. Stealth and persistence are key objectives in backdoor attacks, motivating the design of invisible [nguyen2021wanet, Gao_attack24], semantic [bagdasaryan2021blind, xu2024csba_semantic_atk], optimized [geiping2020witches, sun2024backdoor] and sample-specific [iad, li2021invisible, zhu2025towards] triggers. However, these approaches are limited in FL, because adversaries can only manipulate a limited subset of clients and must ensure their backdoored updates survive the aggregation process to impact the global model. Specifically, FL-specific backdoor attacks can be classified into following two categories.
Single-client Attack. The first backdoor attack method for FL introduced the model replacement attack technic  [bagdasaryan2020backdoor], where a malicious client uploaded a scaled local update to dominate the aggregated model. However, the abnormally large update made it easily detectable under detection [zitengsun01pgd]. To improve stealth and persistence, follow-up works introduced more subtle attack strategies. For example, PGD [zitengsun01pgd] constrained malicious updates within an $\ell_{2}$-ball to evade norm-based detection, while Neurotoxin [Zhengming01Neurotoxin] injected backdoors into parameters associated with inactive neurons to minimize noticeable changes. Other works aimed to further blend backdoor patterns into normal behavior. For example, Edge-case attack [wang2020edgecase] utilized naturally occurring features as triggers, reducing the likelihood of detection, while Chameleon [chameleon] aligned backdoor and benign features using contrastive learning. Beyond stealthiness, a recent work [tdsc25_fba_atk1_durable] analyzes vulnerable layers in NLP models to design a more durable backdoor attack, achieving promising results on LSTM and GPT-2, while remaining limited to specific model architectures and task settings, yet offering valuable insights for general design for durability.
Multi-client Attack. The DBA framework [xie2019dba] leveraged the distributed nature of FL to spread the backdoor across multiple clients, with each client using a distinct local trigger. These independently subtle triggers collectively formed an effective global backdoor, reducing the detectability of any individual client. Expanding on this idea, collusive multi-client attacks [gong2022coordinated_dba, tdsc25_fba_atk2_collusive] have been proposed, in which backdoor triggers are further optimized in a model-aware manner. Such learning-based coordination enables attackers to craft triggers that are both less suspicious and more effective, thereby amplifying the threat posed by distributed poisoning. In contrast, Non-Cooperative Backdoor Attacks (NBA) [nguyen2024nba] involved independent attackers using distinct triggers for different targets, yet still exhibited enhanced stealth comparable to coordinated methods.
@@ 2.3 Backdoor Defenses in Federated Learning
In centralized settings, backdoor defense typically focus on either removing the backdoor from the model [li2021ba_defense_kd, wu2021ba_defense_prune] or filtering trigger samples during training or inference [qiu2021deepsweep, javaheripi2020cleann]. However, these approaches are often impractical in FL due to the defender’s limited capabilities ($i.e.$, the central server): (1) the server lacks access to clean and in-distribution data, making data-dependent removal methods like pruning, retraining, or fine-tuning infeasible. (2) the server cannot control the local training process, rendering sample filtering and robust training strategies infeasible. To address these limitations, FL-specific backdoor detection mechanisms have been developed, generally falling into below two categories.
@@
Backdoor Effect Mitigation methods aim to reduce the impact of backdoor behaviors in the global model by suppressing anomalous neurons across local models. Early works such as Weak-DP [zitengsun01pgd] applied norm clipping and noise injection to weaken abnormal neuron effects. Statistical approaches like geometric median [RFA] and Trimmed-Mean [blanchard2017machine] enforced robustness by aggregating updates in a statistically neutral manner. RLR [ozdayi2021RLR] further enhanced mitigation granularity by assigning parameter-wise adaptive learning rates. FLPurifier [zhang2024flpurifier] achieved stronger backdoor suppression by performing adaptive aggregation at the classifier level under a modified FL protocol that decoupled encoder and classifier training. However, these methods operate all clients equally, often diminishing benign model contributions and harming overall performance under the standard FL protocol. This limitation has led to the rise of the following detection-based methods as a mainstream alternative.
@@
Backdoor Client Detection methods aim to identify and exclude malicious local models during aggregation, typically by quantifying distance between benign and malicious client models to find outliers. For instance, Multi-Krum [multiKrum] leveraged Euclidean distance, while Foolsgold [foolsgold], FLTrust [cao2021fltrust], and ShieldFL [ma2022shieldfl] adopted cosine similarity, all applied directly on full model weights. Rflbat [wang2022rflbat] extended this by computing distances in a top-$k$ PCA-projected space, aiming to reduce noise and improve efficiency, though at the cost of some information loss. Beyond model parameters, DeepSight [rieger2022deepsight] incorporated prediction-level discrepancies of local models, enabling a more comprehensive analysis. Extending beyond pure detection, Flame [nguyen2022flame] combined clustering-based client identification with gradient norm-based mitigation, positioning itself at the intersection of detection and mitigation strategies and demonstrating strong empirical performance. FLGuardian [tifs25_fl_defens_poison], a recent study addressing untagged backdoor attack, performed layer-wise model analysis using similarity metrics and clustering algorithms, reflecting a growing shift toward more fine-grained detection in the parameter space. However, these methods assume malicious models are outliers, making them sensitive to data heterogeneity and ineffective against stealthy attacks.
@@
Since all above detection methods operate without any server-side intervention, we categorize them as passive detection. In contrast, proactive detection explicitly intervenes in FL training to facilitate backdoor identification. The recent pioneering work Indicator [li2024backdoorindicator] adopts this strategy by injecting OOD mappings into the global model, achieving improved robustness under data heterogeneity. However, it still suffers from high false-positive rates due to OOD prediction bias. Recent work [tdsc25_fba_dfs2_proactive_poi] has also applied the proactive detection paradigm to poisoning attacks by amplifying directional deviations between malicious and benign updates through broadcasting a scaled and perturbed, detection-oriented surrogate of aggregated gradients. However, this approach relies on the explicit accuracy-degradation objective of poisoning attacks, which is less stealthy than backdoors and easier to detect, and further depends on parameter-space gradient comparisons that are susceptible to non-i.i.d. bias. In summary, effective proactive backdoor detection in FL remains an open challenge.

### v4

# Coward: Collision-Based OOD Watermarking for Practical Proactive Federated Backdoor Detection

@@ Abstract
Backdoor detection is currently the mainstream defense against backdoor attacks in federated learning (FL), where a small number of malicious clients can upload poisoned updates to compromise the federated global model. Existing backdoor detection techniques fall into two categories, passive and proactive, depending on whether the server proactively intervenes in the training process. However, both of them have practical limitations: passive detection methods are disrupted by common non-i.i.d. data distributions and random participation of FL clients, whereas current proactive detection methods are misled by an inevitable out-of-distribution (OOD) bias because they rely on backdoor coexistence effects. To address these issues, we introduce a novel proactive detection method dubbed Coward, inspired by our discovery of multi-backdoor collision effects, in which consecutively planted, distinct backdoors significantly suppress earlier ones. Correspondingly, we modify the federated global model by injecting a carefully designed backdoor-collided watermark, implemented via regulated dual-mapping learning on OOD data. This design not only enables an inverted detection paradigm compared to existing proactive methods, thereby naturally counteracting the adverse impact of OOD prediction bias, but also introduces a low-disruptive training intervention that inherently limits the strength of OOD bias, leading to significantly fewer misjudgments. Extensive experiments on benchmark datasets show that Coward achieves state-of-the-art performance and effectively alleviates OOD bias.
keywords:
Backdoor Defense , Backdoor Attack , Federated Learning , AI Security , Trustworthy ML
journal: Pattern Recognition
[…]
[tsinghua] organization=Tsinghua University, country=China
[…]
[ecnu] organization=East China Normal University, country=China
[…]
[ntu] organization=Nanyang Technological University, country=Singapore
@@ 1 Introduction
Federated Learning (FL) [25] has emerged as a powerful paradigm for privacy-preserving machine learning, enabling distributed clients to collaboratively train models without sharing raw data, and is becoming a promising basis for trustworthy pattern recognition applications. By keeping data locally owned and processed, FL inherently strengthens privacy and copyright protection, and unlocks the value of proprietary domain-specific and personal data held by private organizations and individuals. This privacy-native property makes FL particularly suitable for building secure and personalized intelligence that goes beyond models trained solely on generic public data [32, 35, 45]. As a result, FL has been extensively studied across many pattern recognition domains, such as healthcare and finance [22, 13, 1].
@@
Figure 1: Our backdoor-collided detection versus the existing backdoor-coexistent method. With our backdoor-collided watermark, we enable an inverted detection paradigm that effectively mitigates misleading effects caused by the OOD prediction bias, a key challenge that limits current proactive detection methods.
@@
However, the privacy-preserving design of federated learning is a double-edged sword. While keeping raw data local effectively protects sensitive information, it also restricts the server’s ability to inspect participant behavior during training. This restricted observability creates an inevitable vulnerability to model-level attacks, where the integrity of the federated global model can be compromised by covertly manipulated local models, most prominently through backdoor attacks [19, 8, 20]. In general, backdoor attacks cause models to behave normally on benign inputs while producing harmful outputs only under specific trigger conditions, rendering the malicious behavior stealthy and difficult to detect. Once the malicious models are aggregated together with benign ones, the backdoor is seamlessly propagated to the global model, thereby manipulating the entire federated system [3, 14].
@@
To detect and exclude malicious backdoor models, passive detection methods have been proposed [34, 27], which do not intervene in the FL process and assume that malicious models manifest observable model-level deviations as outliers. They rely on similarity-based measurements over either model parameters or outputs and identify anomalous models either through direct thresholding or by further applying clustering to expose structurally isolated outliers. However, we reveal that these methods are vulnerable to client non-i.i.d. data distributions and participation randomness, which are typical in FL [25]. These factors lead to highly dynamic and divergent benign local models that may appear as outliers, significantly interfering with detection. We refer to this limitation as the non-i.i.d. divergence challenge, as analyzed in Section 3.2.
@@
Recently, the pioneering work BackdoorIndicator [17] proposed the first federated proactive detection strategy, where the server explicitly intervenes during training to induce malicious models to exhibit behaviors distinct from benign ones. Specifically, BackdoorIndicator injects OOD patterns designed to be easily retained by malicious models while tending to be forgotten by benign ones, such that stronger pattern retention in a local model indicates a malicious client. By shifting the detection basis from model discrepancies to differences in client responses to intervention actions, proactive methods naturally avoid the sensitivity to non-i.i.d. distributions that often undermines passive detection methods. However, due to the privacy constraints of FL that prevent access to in-distribution client data, proactive training interventions at the server are restricted to operating via OOD data. As a result, detection outcomes become susceptible to the OOD prediction bias of deep neural networks (DNNs) [26, 15]. This bias causes benign local models to be easily misidentified as malicious clients. We provide a detailed analysis in Section 3.3 and refer to this limitation as the OOD bias challenge.
@@
To tackle the above challenges, we propose Collision-based OOD WAtermarking for Robust federated backdoor Detection, termed Coward. Our method is inspired by the finding of the multi-backdoor collision effect, where a consecutively planted distinct backdoor significantly interferes with the one previously planted. To this end, we detect backdoor models by evaluating whether the server-injected conflicting global watermark has been erased during local training. Concretely, the global OOD watermark consists of a regulated base OOD mapping and a targeted watermark mapping, where the former assigns a static correspondence from the OOD sample space to the task label space, and the latter connects triggered OOD samples to the watermark target class. In general, our method alleviates the OOD bias challenge through three aspects: 1) Unlike prior approaches, our collision-based detection treats high inspection accuracy as the signal of benign behavior, making it naturally compatible with the OOD bias–induced high-confidence predictions; 2) Our watermark-based detection inspects the presence of the trigger-to-target mapping rather than the OOD sample-to-target mapping, mitigating the direct impact of OOD prediction bias. 3) Our method enforces explicit class-wise label assignments to OOD samples to reduce random prediction fluctuations. We demonstrate (1) and (2) in Figure 1, and verify (3) in Figure 7.
@@
Our main contributions are summarized as follows: (1) We revisit existing detection methods for FL backdoor attacks and identify two key challenges: non-i.i.d. divergence and OOD prediction bias. We show how these challenges undermine existing detection methods, leading to high benign-client misclassification. (2) We find an intriguing phenomenon, dubbed the multi-backdoor collision effect, where a consecutively planted distinct backdoor can significantly suppress the previously implanted one. Based on this finding, we design a simple yet effective proactive backdoor detection method, $i.e.$, Coward, that effectively mitigates OOD bias and achieves more accurate detection. (3) Extensive experiments on benchmark datasets verify the effectiveness of our method and its resistance to potential adaptive backdoor attacks.
@@ 2.1 Federated Learning
Federated learning enables multiple decentralized clients to collaboratively train a shared model without exchanging their private data. Based on how data is partitioned, FL is typically categorized into horizontal FL (HFL), where clients share the same feature space but different samples, and vertical FL (VFL), where clients share samples but differ in input fields [44]. In this paper, we focus on the HFL setting, where a central server coordinates model training by periodically selecting a subset of clients to receive the current global model, update it using local data, and return the updated parameters. The server then aggregates these updates to form a new global model, and this process repeats iteratively, progressively improving model performance while keeping all training data local. Specifically, given the selected client set $\mathcal{S}^{t-1}$ in round t1t-1, the global model $\bm{w}^{t}$ in round $t$ is:
@@
where $N_{k}$ and $N$ denote the data volume of client $k$ and the total data volume across participating clients, respectively. $\Delta$ denotes the model residual with respect to the previous global model. We highlight three key characteristics that distinguish FL from centralized settings: (1) Multi-party Participation: The global model is jointly trained by multiple clients, each of which contributes to the model privately but cannot individually control the training outcome. (2) Non-i.i.d. Client Data [25]: Data distributions can vary significantly across clients, leading to divergent local updates and making the global optimization layer more challenging. (3) Partial Client Participation[25]: In each round, only a subset of clients is randomly selected due to limited bandwidth and dynamic client availability, increasing the uncertainty of training dynamics.
@@ 2.2 Backdoor Attacks in Federated Learning
Backdoor attacks aim to manipulate neural networks to function normally on clean data but misclassify inputs containing attacker-specified triggers [19, 12]. Data poisoning is the most common backdoor injection method, where a subset of training samples is embedded with triggers and mislabeled to a target class [12]. Stealth and persistence are key objectives in backdoor attacks, motivating the design of invisible [28, 10], semantic [42], optimized [36], sample-specific [50] and transformation-aware [20] backdoor strategies. However, these approaches are limited in FL, because adversaries can only control a limited subset of clients and must ensure their backdoored updates survive the aggregation process to impact the global model. Specifically, FL-specific backdoor attacks can be classified into the following categories.
Single-client Attack. The first backdoor attack method for FL introduced the model replacement attack technique [3], where a malicious client uploaded a scaled local update to dominate the aggregated model. However, the abnormally large update made it easily detectable by detection methods [37]. To improve stealth and persistence, follow-up works introduced more subtle attack strategies. For example, PGD [37] constrained malicious updates within an $\ell_{2}$-ball to evade norm-based detection, while Neurotoxin [48] injected backdoors into parameters associated with inactive neurons to minimize noticeable changes. Other works aimed to further blend backdoor patterns into normal behavior. For example, Edge-case attack [38] utilized naturally occurring features as triggers, reducing the likelihood of detection, while Chameleon [8] aligned backdoor and benign features using contrastive learning.
Multi-client Attack. The DBA framework [41] leveraged the distributed nature of FL to spread the backdoor across multiple clients, with each client using a distinct local trigger. These independently subtle triggers collectively formed an effective global backdoor, reducing the detectability of any individual client. Expanding on this idea, collusive multi-client attacks [11, 14] have been proposed, in which backdoor triggers are further optimized in a model-aware manner. Such learning-based coordination enables attackers to craft triggers that are both less suspicious and more effective, thereby amplifying the threat posed by distributed poisoning. In contrast, Non-Cooperative Backdoor Attacks  [29] involved independent attackers using distinct triggers for different targets, yet still exhibited enhanced stealth comparable to coordinated methods.
@@ 2.3 Backdoor Defenses in Federated Learning
In centralized settings, backdoor defenses typically focus on either removing the backdoor from the model [40, 16, 23] or filtering trigger samples during training or inference [33, 46]. However, these approaches are often impractical in FL due to the defender’s limited capabilities ($i.e.$, the central server): (1) the server lacks access to clean and in-distribution data, making data-dependent removal methods like pruning, retraining, or fine-tuning infeasible. (2) the server cannot control the local training process, rendering sample filtering and robust training strategies infeasible. To address these limitations, FL-specific backdoor detection mechanisms have been developed, generally falling into the following two categories.
@@
Backdoor Effect Mitigation methods aim to reduce the impact of backdoor behaviors in the global model by suppressing anomalous neurons across local models. Early works such as Weak-DP [37] applied norm clipping and noise injection to weaken abnormal neuron effects. Statistical approaches like geometric median [31] and Trimmed-Mean [4] enforced robustness by aggregating updates in a statistically neutral manner. RLR [30] further enhanced mitigation granularity by assigning parameter-wise adaptive learning rates. FLPurifier [47] achieved stronger backdoor suppression by performing adaptive aggregation at the classifier level under a modified FL protocol that decoupled encoder and classifier training. However, these methods operate all clients equally, often diminishing benign model contributions and harming overall performance under the standard FL protocol. This limitation has led to the rise of the following detection-based methods as a mainstream alternative.
@@
Backdoor Client Detection methods aim to identify and exclude malicious local models during aggregation, typically by quantifying distance between benign and malicious client models to find outliers. For instance, Multi-Krum [5] leveraged Euclidean distance, while Foolsgold [9], FLTrust [6], and ShieldFL [24] adopted cosine similarity, all applied directly on full model weights. Rflbat [39] extended this by computing distances in a top-$k$ PCA-projected space, aiming to reduce noise and improve efficiency, though at the cost of some information loss. Beyond model parameters, DeepSight [34] incorporated prediction-level discrepancies of local models, enabling a more comprehensive analysis. Extending beyond pure detection, Flame [27] combined clustering-based client identification with gradient norm-based mitigation, positioning itself at the intersection of detection and mitigation strategies and demonstrating strong empirical performance. FLGuardian [49], a recent study addressing untagged backdoor attack, performed layer-wise model analysis using similarity metrics and clustering algorithms, reflecting a growing shift toward more fine-grained detection in the parameter space. However, these methods assume malicious models are outliers, making them sensitive to data heterogeneity and ineffective against stealthy attacks.
@@
Since all the above detection methods operate without any server-side intervention, we categorize them as passive detection. In contrast, proactive detection explicitly intervenes in FL training to facilitate backdoor identification. The recent pioneering work Indicator [17] adopts this strategy by injecting OOD mappings into the global model, achieving improved robustness under data heterogeneity. However, it still suffers from high false-positive rates due to OOD prediction bias. Recent work [43] has also applied the proactive detection paradigm to poisoning attacks by amplifying directional deviations between malicious and benign updates through broadcasting a scaled and perturbed, detection-oriented surrogate of aggregated gradients. However, this approach relies on the explicit accuracy-degradation objective of poisoning attacks, which is less stealthy than backdoors and easier to detect, and further depends on parameter-space gradient comparisons that are susceptible to non-i.i.d. bias. In summary, effective proactive backdoor detection in FL remains an open challenge.